Suppose that the economy is characteised by the following behavioral equations C=Co +c1Yd
Where : Co=280
I=270
G=300
T=200
Marginal propensity to save (MPS) is 0.4(or 40%)
Solve for
a) equlibrum Gdp (Y)
b) disposable income (Yd)
C) Consumption spending (C)
1
Expert's answer
2020-04-23T11:51:53-0400

a) "Y= I+G"

Y=300+270=570

b) "Yd=I-T"

Yd= 270-200= 70

C)"C=Co +c1Yd"

C=280-C1*70